# Gauging or extending bulk and boundary conformal field theories: Application to bulk and domain wall problem in topological matter and their descriptions by (mock) modular covariant

### v2

@@ II Boundary and defect conformal field theories: chiral extension, symmetry charge and amplitude
For simplicity, we restrict our attention to a unitary model with $S^{\alpha}_{0}>0$ for all $\alpha$ where 00 represents the vacuum of CFT. (However, one can apply a similar analysis for a class of nonunitary minimal models by replacing the exact vacuum with the effective vacuum. This exchange corresponds to the Galois shuffle in [162, 163, 73, 12].)
@@
The topological symmetry operator which corresponds to a generalized global symmetry [166] can be written as,
@@
where $m$ and $\overline{m}$ is the label of descendant states and $P$ is the corresponding projection operator. The topological symmetry operator can become noninvertible and this has captured attention recently in the classification of TOs[22]. If one interprets $\mathcal{Q}$ as a symmetry (or integral of motion), the phase and amplitude of $S^{\alpha^{\prime}}_{\alpha}/S^{0}_{\alpha}$ plays a fundamental role. For this purpose, we introduce the following quantity,
@@
where $r$ is a real number with $r\geq 0$ and $Q_{\alpha}(\alpha^{\prime})$ defined by modulo $1$ is the charge of the quantum states labeled by $\alpha^{\prime}$ under the symmetry $\mathcal{Q}_{\alpha}^{\text{mon}}$. For example, when $\mathcal{Q}_{\alpha}$ corresponds to a $Z_{N}$ symmetry, the amplitude produces $r_{\alpha\alpha^{\prime}}=1$ and the phase produces the $Z_{N}$ symmetry charge or anomaly. Hence it is reasonable to name $r_{\alpha\alpha^{\prime}}$ as symmetry amplitude for the quantum state labeled by $\alpha^{\prime}$ under symmetry $\mathcal{Q}_{\alpha}$. Moreover, when the symmetry is noninvertible, some matrix elements of $S$ and the corresponding amplitudes should vanish. In Section III, we study the implication of the noninvertible symmetry from this view.
@@ II.1 Algebraic data of Ising and Majorana CFT: Bulk and chiral
where the identification from the ring isomorphism is $\{I,\psi,\sigma\}\sim\{I,\epsilon,\sigma_{\text{Bulk}}\}$. $I$ is the identity operator, $\epsilon$ is the energy and $\sigma_{\text{Bulk}}$ is the $Z_{2}$ order operator operator. For later discussions, we stress that we do not assume the chiral and antichiral decomposition without respective cautions, because this decomposition sometimes contains subtelities. Moreover, in general, this is not a tensor decomposition (in category theory, this is called Deligne product [174] and this structure is verified by “abstract nonsense"[175, 2]).
@@
Because this is an extension (or gauging), the number of bulk fields becomes twice. Roughly speaking this extension has been achieved by the introduction of the chiral (antichiral) $Z_{2}$ field $\psi$ ($\overline{\psi}$) and its multiplication to the Ising SFC. The most nontrivial point of this extension is the introduction of the disorder fields $\mu_{\text{Bulk}}$ represented as $\mu_{\text{Bulk}}=\psi\times\sigma_{\text{Bulk}}$, because the naive tensor product representation $\sigma\otimes\overline{\sigma}$ results in the relation $\psi\times(\sigma\otimes\overline{\sigma})=\sigma\otimes\overline{\sigma}$.
@@
After the $Z_{2}$ extension, one can define $Z_{2}$ grading for each particle in the CCFT. For example, one can take $I$ and $e$ as 00, and $\psi$ and $m$ as $1$. One can check its consistency by replacing each object in the double semion algebra to 00 or $1$ and $\times$ to $+$,
@@ III Ordering by bulk relevant perturbation and its interpretation in BCFT
[…]
@@
The conjecture by Cardy in [84] is simple and general (The correspondence between massive RG and BCFT has first appeared in the study of integrable models in our knowledge[96, 97, 98]). The conjectue says that the ground state of $H_{\text{Lat}}$, $|GS\rangle,$ can be approximated by smeared Cardy states $\sum_{\alpha}e^{-\tau_{\alpha}H_{\text{CFT}}}|\alpha\rangle$ constructed from Cardy states $\{|\alpha\rangle\}$. The parameter $\tau_{\alpha}$ is a model parameter that depends on the finite size scaling analysis. However, as we will demonstrate, the contribution from $\tau_{\alpha}$ is not relevant to our purpose and we dropped off the contribution for simplicity. For the readers interested in more detailed calculations, we note several recent works[87, 132, 190, 130],
@@
[…]
@@
Figure 3: Mass-condensation or ordering in BCFT and RG. In a CFT without boundary, the vacuum expectation value of a bulk primary field $\Phi_{\alpha}$ should vanish. Under the operator-state correspondence, this vanishing corresponds to the vanishing of the expectation value of LSM twist operator $U_{\alpha}$ corresponding to the many-body elementary excitation. Under RG, the ground state expectation value of the LSM operator becomes finite and this is a mass-condensation in the QFT side. The same condensation can be mimicked by introducing boundary to the same CFT, and this represents (massive) QFT/BCFT correspondence summarized by Cardy[84].
@@
Comparing Eq. (29) and Eq. (31), one can map the problem of bulk RG flow to the classification of boundary conditions in BCFT and this is a concise representation of the proposal by Cardy [84] (Fig. 3). In other words, the vanishing of the element of the modular $S$ matrix restricts RG flows or the type of gapped phase determined by the mass condensation. The appearance of zero in the modular $S$ matrix elements implies obstruction for the condensation. Moreover, we note that the appearance of zero in the elements of a modular $S$ matrix results in the appearance of a noninvertible symmetry because of the matrix representation of the Verlinde line or Eq. (8) as a topological symmetry operator. Hence one can obtain the following correspondence,
[…]
@@
[…]
If $r_{\alpha\beta}=0$ the system cannot flow to the gapped phase $|\beta\rangle\rangle$ under the perturbation $\Phi_{\alpha}$(and its Hermitian conjugate). More generally, the other perturbations inducing $\langle GS|\Phi_{\alpha}|GS\rangle\neq 0$ cannot flow the system to the gapped phase $|\beta\rangle\rangle$.
@@
[…]
Hence one can state that the zero of the symmetry amplitude can restrict the phase diagram only by assuming RG flows from CFTs. This might be thought of as a noninvertible analog of symmetry-protected RG flow, but we have concentrated our attention on the symmetry amplitude, not a charge. If one induces the noninvertible symmetry $\mathcal{Q}_{\alpha}$ with $r_{\alpha\beta}=0$ as the integral of motion in the usual sense, this prohibits the existence of the perturbation $\Phi_{\beta}$. However, it should be mentioned that the vanishing of the amplitude $r_{\alpha\beta}=0$ does not necessarily mean $\mathcal{Q}_{\alpha}\Phi_{\beta}=0$ in considering lattice analog, because $\mathcal{Q}_{\alpha}\Phi_{\beta}$ can be a disorder operator. Because a disorder operator is out of the scope of the local operators in the original bosonic CFT, the amplitude becomes zero. Inversely, the vanishing of the amplitude may imply the existence of a disorder operator, but this may require more careful arguments on the object $\mathcal{Q}_{\alpha}\Phi_{\beta}$. Hence, the following conjectural phenomenology will be important for future studies,
[…]
@@
Considering the Lorentz invariance of the theory, it is useful to study the relationship between $\{\mathcal{Q}_{\alpha}\Phi\}$ and twisted theory by $\{\mathcal{D}_{\alpha}\Phi\}$ where $\mathcal{D}_{\alpha}$ is a topological defect or domain wall (Fig.4). The action of $\mathcal{D}_{\alpha}$ to the bulk fields $\Phi$ can be implemented by the chiral and antichiral decomposition of $\Phi$, roughly speaking. Hence by applying the modular $S$ transformation to the bulk fields, one can obtain the correspondence as follows.
@@
Figure 4: The connection between symmetry and defect under modular $S$ transformation. The bondary states $|\beta\rangle\rangle$ can also be considered as defect condensation $\mathcal{Q}_{\beta}|0\rangle\rangle$[206, 208]. Hence, in a large class of models, the combination of defects and symmetry governs the RG to gapped phase.
We note here the application of analysis in this section to the Ising CFT and discuss its interpretation in the quantum spin chain. There exist three chiral primary fields $\{I,\psi,\sigma\}$, and one can label the corresponding Cardy states by them. $I$ and $\psi$ exchanges each other, and they correspond to ordered phase breaking $Z_{2}$ symmetry. The perturbation that triggers this flow is the primary fields $\sigma_{\text{Bulk}}$ and the $Z_{2}$ spin flip changing the sign of the coupling constant exchange each phase. The boundary state $|\sigma\rangle\rangle$ has to correspond to the invariant phase under the $Z_{2}$ spin flip. Hence this corresponds to the disordered phase. We remark that the $Z_{2}$ invariance of $\sigma$ plays a significant role in this identification.
@@ III.1 Anomaly-free theory and smeared BCFT: Boundary states outside of the boundary of lattice models
where $i$ is a label of modular noninvariant sectors, $a$ is that of $Z_{N}$ invariant sectors and $p$ is the $Z_{N}$ parity or $Z_{N}$ particle number couting taking $p=0,1,..,N-1$. The $Z_{N}$ simple current is a chiral primary field acting like a generator of $Z_{N}$ symmetry, such as $J\times\phi_{\alpha}=\phi_{\beta}$ where $\phi$ is a primary field. This kind of modular invariants appears in $\{SU(N)_{1}\}^{k}$ WZW models or general $Z_{N}$ symmetric models with $N$ prime number (Similar arguments applies to more general cases by considering the prime decomposition of the number $N$ corresponding to subgroup of $Z_{N}$, and the author will present more general results elsewhere). We have followed the notation in the author’s previous works[12, 32]. The anomaly-free condition[38, 39, 40, 41] can be summarized as follows,
  • ($N$ integer) $h_{J^{k}}=$integer for all $k$ and $N$,
  • ($N$ even case) $h_{J^{k}}=$ integer for even $k$ and $h_{J^{k}}=$ half-integer for odd $k$,
@@
Because of the extension, the index $a$ in Eq. (38) is generalized to $(a,p)$. More mathematically, the simple current extension of a bosonic CCFT is an extension of $Z_{N}$ symmetric MTC to $Z_{N}$ graded chiral algebra (In algebraic level, the theory $\{\phi\}$ is well-defined, but it is not sure this is in the scope of the existing category theory. This might correspond to the premodular fusion category (PMF) or superfusion category, but this identification is not conclusive. We thank Ingo Runkel and Steve Simon for reminding us the significance of this aspect.).
@@
Figure 5: Existence of a disordered state and its implication to an anomaly-free model.
These boundary conditions labeled by $i$ and $a$ correspond to the boundary conditions of a quantum spin chain or its gauged version because the fusion rule produces only positive amplitude. Hence this representation is called nonnegative-integer-matrix (NIM) representation. However, in the $Z_{N}$ gauged theories, there can exist twisted boundary conditions,
@@
where $\omega=e^{\frac{2\pi i}{N}}$ is the $N$-th root of unity and $l=0,..,N-1$ is the label of twist. One can implement this twist by applying the $Z_{N}$ parity operator. Hence the set of boundary states is labeled by $(i,l)$ and $(a,l)$. The most significant property of this representation is,
@@
under the application $\mathcal{Q}_{\sigma}$ corresponding to the KW dual transformation[235, 236, 237]. Hence the duality is a two-to-one mapping and noninveritible in bosonic representation.
@@
Hence one can confirm that the fermionic theory is the correct extension of bosonic one, by taking the subset $\{I,\psi,\sigma\}$.
The states $|e\rangle\rangle\pm|m\rangle\rangle$ correspond to the fermionic zero modes in the Ramond sector and also to $Z_{2}$ extension of disordered phase with a unique ground state. Moreover, under assuming some additional symmetries from the lattice, the disordered phase corresponds to the Haldane phase. There exists a general correspondence between the charged fermionic zero modes and the Haldane phase of spin $S$ antiferromagnetic Heisenberg chain with $S$ odd integer by studying the corresponding $SU(2)_{2S}$ Wess-Zumino-Witten (WZW) model[238, 239]. The $Z_{2}$ symmetric operator labeled by $a$ becomes $Z_{2}$ charged objects, only when the conformal dimension of the $Z_{2}$ simple current $J$ is half-integer. By considering the corresponding smeared Cardy states labeled by $a$, this explains the existing classification in [183, 184, 185, 186] in a simple but reasonable way. The conformal dimension of the $Z_{2}$ simple current in the corresponding WZW model becomes a half-integer only when the spin $S$ is an odd integer. The mapping between the Haldane and fermionic topological phase in [240] corresponds to this observation. Moreover, this argument seems to be connected to the degeneracies of the entanglement spectrum because of its close connection to zero modes $(a,p)$. Hence one can obtain the following phenomenology,
@@
In other words, the underlying bosonic or fermionic statistics [182] of $Z_{N}$ particle $J$ govern the SPT classifications. This classification will be useful for the $Z_{2N}$ models because these models addmit both fermionic and bosonic statistics. Similar periodicity can be seen in [241] and a possible generalization to massless RG flows can be seen in [242].

### v3

@@ II Boundary and defect conformal field theories: chiral extension, symmetry charge and amplitude
For simplicity, we restrict our attention to a unitary model with $S^{\alpha}_{0}>0$ for all $\alpha$ where 0 represents the vacuum of CFT. (However, one can apply a similar analysis for a class of nonunitary minimal models by replacing the exact vacuum with the effective vacuum. This exchange corresponds to the Galois shuffle in [163, 164, 73, 12].)
@@
From the arguments in [158] the topological symmetry operator (or Verlinde line), which can be interpreted as a proto type of generalized global symmetry [167, 168], can be written as,
@@
where $m$ and $\overline{m}$ is the label of descendant states and $P$ is the corresponding projection operator. Because these operators are written as a summation of projections, they commute with the Hamiltonian and behave as symmetries (or conserved quantities). The topological symmetry operator can become noninvertible and this has captured attention recently in the classification of TOs[22]. If one interprets $\mathcal{Q}$ as a symmetry (or integral of motion), the phase and amplitude of $S^{\alpha^{\prime}}_{\alpha}/S^{0}_{\alpha}$ play a fundamental role. For this purpose, we introduce the following quantity,
@@
where $r$ is a real number with $r\geq 0$ and $Q_{\alpha}(\alpha^{\prime})$ defined by modulo $1$ is the charge of the quantum states labeled by $\alpha^{\prime}$ under the symmetry $\mathcal{Q}_{\alpha}^{\text{mon}}$. For example, when $\mathcal{Q}_{\alpha}$ corresponds to a $Z_{N}$ symmetry, the amplitude produces $r_{\alpha\alpha^{\prime}}=1$ and the phase produces the $Z_{N}$ symmetry charge or anomaly. Hence it is reasonable to name $r_{\alpha\alpha^{\prime}}$ as symmetry amplitude for the quantum state labeled by $\alpha^{\prime}$ under symmetry $\mathcal{Q}_{\alpha}$. Moreover, when the symmetry is noninvertible, some matrix elements of $S$ and the corresponding amplitudes should vanish. In Section III, we study the implications of the noninvertible symmetry from this view.
@@ II.1 Algebraic data of Ising and Majorana CFT: Bulk and chiral
where the identification from the ring isomorphism is $\{I,\psi,\sigma\}\sim\{I,\epsilon,\sigma_{\text{Bulk}}\}$. $I$ is the identity operator, $\epsilon$ is the energy and $\sigma_{\text{Bulk}}$ is the $Z_{2}$ order operator operator. For later discussions, we stress that we do not assume the chiral and antichiral decomposition without respective cautions, because this decomposition sometimes contains subtelities. Moreover, in general, this is not a tensor decomposition (in category theory, this is called Deligne product [176] and this structure is verified by “abstract nonsense"[177, 2]). In any case, the above three algebraic objects reproduce the established bosonic modular invariant $|\chi_{I}|^{2}+|\chi_{\psi}|^{2}+|\chi_{\sigma}|^{2}$.
@@
Because this is an extension (or gauging), the number of bulk fields becomes twice. Roughly speaking this extension has been achieved by the introduction of the chiral (antichiral) $Z_{2}$ field $\psi$ ($\overline{\psi}$) and its multiplication to the Ising SFC. The most nontrivial point of this extension is the introduction of the disorder fields $\mu_{\text{Bulk}}$ represented as $\mu_{\text{Bulk}}=\psi\times\sigma_{\text{Bulk}}$, because the naive tensor product representation $\sigma\otimes\overline{\sigma}$ results in the relation $\psi\times(\sigma\otimes\overline{\sigma})=\sigma\otimes\overline{\sigma}$. Because of the extension, only the Neveu-Schwarz sector $\{I,\psi,\overline{\psi},\epsilon\}$ forms the modular $S$ invariant $|\chi_{I}+\chi_{\psi}|^{2}$ and the locality is reduced to modular $T^{2}$ invariance. The other sector, $\{\sigma_{\text{Bulk}},\mu_{\text{Bulk}}\}$ known as the Ramond sector only forms modular covariant $2|\chi_{\sigma}|^{2}$.
@@
After the $Z_{2}$ extension, one can define $Z_{2}$ grading for each particle in the CCFT. For example, one can take II and $e$ as 0, and $\psi$ and $m$ as $1$. One can check its consistency by replacing each object in the double semion algebra to 0 or $1$ and $\times$ to ++,
@@ III Ordering by bulk relevant perturbation and its interpretation in BCFT
$\{\text{CFT/TQFT}\}\overset{\text{RG}/\text{Gluing}}{\Rightarrow}\{\text{BCFT/CCFT}\}$ (27)
@@
The conjecture by Cardy in [84] is simple and general (The correspondence between massive RG and BCFT has first appeared in the study of integrable models in our knowledge[96, 97, 98]). The conjectue says that the ground state of $H_{\text{Lat}}$, $|GS\rangle,$ can be approximated by smeared Cardy states $\sum_{\alpha}e^{-\tau_{\alpha}H_{\text{CFT}}}|\alpha\rangle$ constructed from Cardy states $\{|\alpha\rangle\}$. The parameter $\tau_{\alpha}$ is a model parameter that depends on the finite size scaling analysis. However, as we will demonstrate, the contribution from $\tau_{\alpha}$ is not relevant to our purpose and we dropped the contribution for simplicity. For the readers interested in more detailed calculations, we note several recent works[87, 133, 192, 131],
@@
$\{\langle 0|\Phi_{\alpha}|0\rangle=0\}_{\text{UV}}\overset{\text{RG}}{\Rightarrow}\{\langle GS|U_{\alpha}|GS\rangle\neq 0\}_{\text{IR}}$ (29)
@@
Figure 3: Mass-condensation or ordering in BCFT and RG. In a CFT without boundary, the vacuum expectation value of a bulk primary field $\Phi_{\alpha}$ should vanish. Under the operator-state correspondence, this vanishing corresponds to the vanishing of the expectation value of the LSM twist operator $U_{\alpha}$ corresponding to the many-body elementary excitation. Under RG, the ground state expectation value of the LSM operator becomes finite and this is a mass-condensation in the QFT side. The same condensation can be mimicked by introducing boundary to the same CFT, and this represents (massive) QFT/BCFT correspondence summarized by Cardy[84].
@@
Comparing Eq. (29) and Eq. (31), one can map the problem of bulk RG flow to the classification of boundary conditions in BCFT and this is a concise representation of the proposal by Cardy [84] (Fig. 3). In other words, the vanishing of the element of the modular $S$ matrix restricts RG flows or the type of gapped phase determined by the mass condensation. The appearance of zero in the modular $S$ matrix elements implies obstruction for the condensation. Moreover, we note that the appearance of zero in the elements of a modular $S$ matrix results in the appearance of a noninvertible symmetry because of the matrix representation of the Verlinde line or Eq. (8) as a topological symmetry operator. Hence, one can obtain the following correspondence,
$\text{Noninvertible symmetry}\sim\text{Obstruction to condensation}.$ (33)
@@
We study a more specific situation. Let us assume the form of the perturbation as $H_{p}=(\lambda/2\pi)\int dx(\Phi_{\alpha}+\Phi_{\alpha}^{\dagger})$ or $H_{p}=(i\lambda/2\pi)\int dx(\Phi_{\alpha}-\Phi_{\alpha}^{\dagger})$ with $S^{\alpha}_{\beta}=0$, where $x$ is the space cordinate and $\lambda$ is a coupling constant and $\dagger$ represents the Hermitian conjugate. In this setting, if the perturbation induces the fixing $\langle GS|\Phi|GS\rangle\neq 0$, this prevents the flow $|GS\rangle\rightarrow|\beta\rangle\rangle$ by contradiction. By identifying the smeared Cardy state $|\beta\rangle$ as a ground state of a gapped phase, we can obtain the following statement;
If $r_{\alpha\beta}=0$ the system cannot flow to the gapped phase $|\beta\rangle\rangle$ under the perturbation $\Phi_{\alpha}$(and its Hermitian conjugate). More generally, the other perturbations inducing $\langle GS|\Phi_{\alpha}|GS\rangle\neq 0$ cannot flow the system to the gapped phase $|\beta\rangle\rangle$. For the readers intersted in nontrivial relations between noninvertible symmetry and RG flows, we note another recent work by the author[206].
@@
$\text{Noninvertible symmetry}\sim\text{Obstruction to RG flow}.$ (35)
Hence, one can state that the zero of the symmetry amplitude can restrict the phase diagram only by assuming RG flows from CFTs. This might be thought of as a noninvertible analog of symmetry-protected RG flow, but we have concentrated our attention on the symmetry amplitude, not a charge. If one induces the noninvertible symmetry $\mathcal{Q}_{\alpha}$ with $r_{\alpha\beta}=0$ as the integral of motion in the usual sense, this prohibits the existence of the perturbation $\Phi_{\beta}$. However, it should be mentioned that the vanishing of the amplitude $r_{\alpha\beta}=0$ does not necessarily mean $\mathcal{Q}_{\alpha}\Phi_{\beta}=0$ in considering lattice analog, because $\mathcal{Q}_{\alpha}\Phi_{\beta}$ can be a disorder operator by considering the commutation relation between $\mathcal{Q}_{\alpha}$ and $\Phi_{\beta}$(See [83, 207]). Because a disorder operator is out of the scope of the local operators in the original bosonic CFT, the amplitude becomes zero. Inversely, the vanishing of the amplitude may imply the existence of a disorder operator, but this may require more careful arguments on the object $\mathcal{Q}_{\alpha}\Phi_{\beta}$. Hence, the following conjectural phenomenology will be important for future studies,
$\text{Existence of a disorder field}\ \Rightarrow\text{Obstruction to RG flow}.$ (36)
@@
Considering the Lorentz invariance of the theory, it is useful to study the relationship between $\{\mathcal{Q}_{\alpha}\Phi\}$ and twisted theory by $\{\mathcal{D}_{\alpha}\Phi\}$ where $\mathcal{D}_{\alpha}$ is a topological defect or domain wall (Fig.4). The action of $\mathcal{D}_{\alpha}$ to the bulk fields $\Phi$ can be implemented by the chiral and antichiral decomposition of $\Phi$, roughly speaking. Hence, by applying the modular $S$ transformation to the bulk fields, one can obtain the correspondence as follows.
@@
Figure 4: The connection between symmetry and defect under modular $S$ transformation. The bondary states $|\beta\rangle\rangle$ can also be considered as defect condensation $\mathcal{Q}_{\beta}|0\rangle\rangle$[210, 212]. Hence, in a large class of models, the combination of defects and symmetry governs the RG to the gapped phase.
We note here the application of analysis in this section to the Ising CFT and discuss its interpretation in the quantum spin chain. There exist three chiral primary fields $\{I,\psi,\sigma\}$, and one can label the corresponding Cardy states by them. II and $\psi$ exchange each other, and they correspond to the ordered phase breaking $Z_{2}$ symmetry. The perturbation that triggers this flow is the primary fields $\sigma_{\text{Bulk}}$ and the $Z_{2}$ spin flip changing the sign of the coupling constant exchange each phase. The boundary state $|\sigma\rangle\rangle$ has to correspond to the invariant phase under the $Z_{2}$ spin flip. Hence this corresponds to the disordered phase. We remark that the $Z_{2}$ invariance of $\sigma$ plays a significant role in this identification.
@@ III.1 Anomaly-free theory and smeared BCFT: Boundary states outside of the boundary of lattice models
where $i$ is a label of modular noninvariant sectors, $a$ is that of $Z_{N}$ invariant sectors and $p$ is the $Z_{N}$ parity or $Z_{N}$ particle number couting taking p=0,1,..,N1p=0,1,..,N-1. The $Z_{N}$ simple current is a chiral primary field acting like a generator of $Z_{N}$ symmetry, such as $J\times\phi_{\alpha}=\phi_{\beta}$ where $\phi$ is a primary field. This kind of modular invariants appears in $\{SU(N)_{1}\}^{k}$ WZW models or general $Z_{N}$ symmetric models with $N$ prime number (Similar arguments applies to more general cases by considering the prime decomposition of the number $N$ corresponding to subgroup of $Z_{N}$, and the author will present more general results elsewhere). We have followed the notation in the author’s previous works[12, 32]. The anomaly-free condition[38, 39, 40, 41] can be summarized as follows,
  • (NN integer) $h_{J^{k}}=$integer for all $k$ and $N$,
  • (NN even case) $h_{J^{k}}=$ integer for even $k$ and $h_{J^{k}}=$ half-integer for odd $k$,
@@
Because of the extension, the index $a$ in Eq. (38) is generalized to $(a,p)$. More mathematically, the simple current extension of a bosonic CCFT is an extension of $Z_{N}$ symmetric MTC to $Z_{N}$ graded chiral algebra (In algebraic level, the theory $\{\phi\}$ is well-defined, but it is not sure this is in the scope of the existing category theory. This might correspond to the premodular fusion category (PMF) or superfusion category, but this identification is not conclusive. We thank Ingo Runkel and Steven Simon for reminding us the significance of this aspect.).
@@
Figure 5: Existence of a disordered state and its implication to an anomaly-free model. If we assume an $Z_{N}$ symmetric anomaly-free theory flows to a nondegenerate disordered phase labelled by $a$, the anomaly freeness imposes the existence of zero modes $(a,p)$. These zero modes only appear when applying the $Z_{N}$ extension to the resultant gapped system and are absent in the original spin or bosonic system.
These boundary conditions labeled by $i$ and $a$ correspond to the boundary conditions of a quantum spin chain or its gauged version because the fusion rule produces only positive amplitudes. Hence this representation is called nonnegative-integer-matrix (NIM) representation. However, in the $Z_{N}$ gauged theories, there can exist twisted boundary conditions,
@@
where $\omega=e^{\frac{2\pi i}{N}}$ is the $N$-th root of unity and l=0,..,N1l=0,..,N-1 is the label of twist. One can implement this twist by applying the $Z_{N}$ parity operator. Hence, the set of boundary states is labeled by $(i,l)$ and $(a,l)$. The most significant property of this representation is,
@@
under the application $\mathcal{Q}_{\sigma}$ corresponding to the KW dual transformation[239, 240, 241]. Hence, the duality is a two-to-one mapping and noninvertible in bosonic representation.
@@
Hence, one can confirm that the fermionic theory is the correct extension of the bosonic one, by taking the subset $\{I,\psi,\sigma\}$.
The states $|e\rangle\rangle\pm|m\rangle\rangle$ correspond to the fermionic zero modes in the Ramond sector and also to the $Z_{2}$ extension of the disordered phase with a unique ground state. Moreover, under assuming some additional symmetries from the lattice, the disordered phase corresponds to the Haldane phase. There exists a general correspondence between the charged fermionic zero modes and the Haldane phase of the spin $S$ antiferromagnetic Heisenberg chain with $S$ odd integer by studying the corresponding $SU(2)_{2S}$ Wess-Zumino-Witten (WZW) model[242, 243]. The $Z_{2}$ symmetric operator labeled by $a$ becomes $Z_{2}$ charged objects, only when the conformal dimension of the $Z_{2}$ simple current $J$ is half-integer. By considering the corresponding smeared Cardy states labeled by $a$, this explains the existing classification in [185, 186, 187, 188] in a simple but reasonable way. The conformal dimension of the $Z_{2}$ simple current in the corresponding WZW model becomes a half-integer only when the spin $S$ is an odd integer. The mapping between the Haldane and fermionic topological phase in [244] corresponds to this observation. Moreover, this argument seems to be connected to the degeneracies of the entanglement spectrum because of its close connection to zero modes $(a,p)$. Hence, one can obtain the following phenomenology,
@@
In other words, the underlying bosonic or fermionic statistics [184] of $Z_{N}$ particle $J$ govern the SPT classifications. This classification will be useful for the $Z_{2N}$ models because these models admit both fermionic and bosonic statistics. Similar periodicity can be seen in [245] and a possible generalization to massless RG flows can be seen in [246].